Diffusion Morning Report 11/19/10

Ted Berg tberg at stanford.edu
Fri Nov 19 06:59:58 PST 2010


Good Morning All,
     looking good this morning. Epi had lift issues. Tylan 1 possible 
temp issues and ebeam still having computer issues. Burn box cleaned 
this morning. No other issues at this time.

> 					
> 	Date: 	11/19/10 	THICKNESS-Change at 		Priority
> Tube Name/Number 	Status 				
> Tylan 1 	*yellow* 	*Possible temp fluctuations* 	** 	** 	*2*
> Tylan 2 	*up* 	** 	** 	** 	**
> Tylan 3 	*up* 	** 	** 	** 	**
> Tylan 4 	*up* 	** 	** 	** 	**
> Tylan 5-BBr3 	*up* 	** 	** 	** 	**
> Tylan 6-POCl3 	*up* 	** 	** 	** 	**
> Tylan 7-TEOS2 	*up* 	** 	*.25/30* 	** 	**
> Tylan 8-SiGe 	*up* 	** 	*20.1/30* 	** 	**
> Tylan 9-Poly 	*up* 	** 	*12.9/30* 	** 	**
> Tylan 10-Nitride 	*up* 	** 	*16.86/30* 	** 	**
> Tylan 11-LTO 	*up* 	** 	*2.5/30* 	** 	**
> Tylan 12-FGA 	*up* 	*
> * 	** 	** 	**
> Thermco 1 	*up* 	** 	** 	** 	**
> Thermco 2 	*up* 	** 	** 	** 	**
> Thermco 3 	*future* 	** 	** 	** 	**
> Thermco4 	*future* 	** 	** 	** 	**
> Thermco Poly1 	*up* 	** 	** 	** 	**
> Thermco Poly2 	*up* 	** 	** 	** 	**
> Thermco Nitride1 	*up* 	** 	** 	** 	**
> Thermco bpsg 	*yellow* 	*testing* 	** 	** 	**
> AMAT Epi 	*yellow* 	*testing and running* 	** 	** 	*1*
> TEL Spa 	*up* 	** 	** 	** 	**
> Burn Box 	*up* 	** 	** 	** 	**
> Ebeam 	*down* 	*computer issues* 	** 	** 	*3*
> ZYGO 	*up* 	*
> * 	** 	** 	**
> FGA2 	*up* 	** 	** 	** 	**
> SiH4 for furnaces 	** 	*A on line 630 PSI* 	*changeover working 
> well* 	** 	**
>

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20101119/1385f1eb/attachment.html>


More information about the diffusion mailing list